OpenStudy (anonymous):

Simplify the expression. cos^2 x + sin^2 x/cot^2 x - csc^2 x

OpenStudy (anonymous):

Do you mean this : \[frac{cos^2( x) + sin^2 (x)}{cot^2 (x) - cosec^2 (x)}\]

OpenStudy (anonymous):

No. I mean this: \[ \frac{ \cos ^{2} x + \sin ^{2} x}{ \cot ^{2} x - \csc ^{2} x }\]

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

I am seeing Math Processing Error.. Just wait I am refreshing my page so that Latex can upload properly..

OpenStudy (anonymous):

Okay :)

OpenStudy (anonymous):

Yeah I was almost there.. :)

OpenStudy (anonymous):

Do you know about Trigonometric Identities??

OpenStudy (anonymous):

Kind of

OpenStudy (anonymous):

I mean you must remember these: \[1. \quad \sin^2(x) + \cos^2(x) = 1\] \[2. \quad 1 + \tan^2(x) = \sec^2(x)\] \[3. \quad 1 + \cot^2(x) = cosec^2(x)\]

OpenStudy (anonymous):

Remember??

OpenStudy (anonymous):

Sooo it'd be 1?

OpenStudy (anonymous):

No wait, would it be -1 because I'm dividing?

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

Numerator will give you 1 but denominator will give you -1.. So 1 divided by -1 will be how much??

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

Excellent... :)

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Welcome dear..